OpenStudy (anonymous):

will medal and fan for help (so help please) find the 50th term of the sequence 5, -2, -9, -16, ... A. -352 B. -343 C. -338 D. -331

OpenStudy (er.mohd.amir):

here common diff d=-2-5=-7 and a=5 n th term=a+(n-1)d

OpenStudy (er.mohd.amir):

where n is number of term a is first term and d is common difference

OpenStudy (anonymous):

so it would be a(50)=5+(50-1)*-7

OpenStudy (anonymous):

right

OpenStudy (er.mohd.amir):

yes

OpenStudy (anonymous):

which would turn into a(50)=55-1*-7 which would than be a(50)=54*-7 which would produce a(50)=-378

OpenStudy (anonymous):

correct

OpenStudy (er.mohd.amir):

5+(50-1)*-7 =5-7*49

OpenStudy (anonymous):

and then it would be5-343 which would then be -348

OpenStudy (anonymous):

right

OpenStudy (er.mohd.amir):

5-343=-338

OpenStudy (anonymous):

so it would be C. correct

OpenStudy (er.mohd.amir):

yes

OpenStudy (anonymous):

what about another question

OpenStudy (er.mohd.amir):

which one

OpenStudy (anonymous):

find the arithmetic mean of an of a\[a _{n-1}=3.9\] and \[a _{n+1}=7.1\]

OpenStudy (er.mohd.amir):

put 3.9=a+(n-1-1)d and 7.1=a+(n+a-1)d and solve then get a and d

OpenStudy (anonymous):

so it would be 1.7

OpenStudy (anonymous):

i mean 1.6

OpenStudy (er.mohd.amir):

3.9=a+(n-1-1)d 7.1=a+(n+1-1)d subtract it 7.1-3.9=a+nd-a-nd+2d 3.2=2d d=1.6

OpenStudy (anonymous):

so i was right

OpenStudy (er.mohd.amir):

now solve a

OpenStudy (anonymous):

3.6

OpenStudy (er.mohd.amir):

u try to find AM so u can simple add and divide by total numbers of terms

OpenStudy (anonymous):

so i am right